What is the probability of flipping a coin six times in a row and getting three heads and three tails?
Answer #1

Here, n = 6, p = 0.5, (1 - p) = 0.5 and x = 3
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 3)
P(X = 3) = 6C3 * 0.5^3 * 0.5^3
P(X = 3) = 0.3125